### v1

# {textblock}

Abstract
We present AIOS 1.0, a novel platform designed to advance computer-use agent (CUA) capabilities through environmental contextualization. While existing approaches primarily focus on building more powerful agent frameworks or enhancing agent models, we identify a fundamental limitation: the semantic disconnect between how language models understand the world and how computer interfaces are structured. AIOS 1.0 addresses this challenge by transforming computers into contextual environments that language models can natively comprehend, implementing a Model Context Protocol (MCP) server architecture to abstract computer states and actions. This approach effectively decouples interface complexity from decision complexity, enabling agents to reason more effectively about computing environments. To demonstrate our platform’s effectiveness, we introduce LiteCUA, a lightweight computer-use agent built on AIOS 1.0 that achieves a 14.66% success rate on the OSWorld benchmark, outperforming several specialized agent frameworks despite its simple architecture. Our results suggest that contextualizing computer environments for language models represents a promising direction for developing more capable computer-use agents and advancing toward AGI that can interact with digital systems. The source code of LiteCUA is available at GitHub1 and it is also integrated into the AIOS main branch as part of AIOS2.
[Uncaptioned image] 1LiteCUA Code Repository: https://github.com/agiresearch/LiteCUA
[Uncaptioned image] 2AIOS Code Repository: https://github.com/agiresearch/AIOS
1 Introduction
Large language models (LLMs) (Achiam et al., 2023, Team et al., 2023, Liu et al., 2024, Dubey et al., 2024, Yang et al., 2024a) have achieved impressive leaps in understanding knowledge and reasoning (Wei et al., 2022, Yao et al., 2023a, Besta et al., 2023, Jin et al., 2024b, 2025, Chen et al., 2022, Jin et al., 2024a), fuelling a new generation of LLM-powered agents that plan, observe, and act to interact with digital environments (Shen et al., 2023, Wu et al., 2023, Ge et al., 2023, Yao et al., 2023b). Research demonstrates that, when coupled with suitable prompting or fine-tuning, LLM-based agents can effectively plan to fulfill complex tool calling (Patil et al., 2024, Qin et al., 2024, Xu et al., 2023) and programming (Wang et al., 2024, Yang et al., , Zhang et al., 2024) tasks.
Moving beyond calling external tools or executing generated code, researchers have begun exploring whether LLM agents can perform more complex computer-use tasks, leveraging their emerging planning and action capabilities. This exploration has catalyzed the development of benchmarks assessing LLM-based agents in computer-use scenarios across operating systems (Xie et al., 2024, Bonatti et al., 2024), mobile devices (Deng et al., 2024, Rawles et al., 2024), and browsers (Yao et al., 2022, Zhou et al., 2023, Deng et al., 2023, Koh et al., 2024, Pan et al., 2024). Parallel efforts have focused on leveraging LLMs and vision-language models (VLMs) to perceive computer/mobile environments and execute actions such as clicking, typing, and dragging interactive elements to complete tasks. Current approaches to computer-use agents (CUAs) primarily follow two trajectories: (1) developing agent frameworks with specialized modules for critical subtasks like grounding (Agashe et al., 2025, Gou et al., 2025), search (Yu et al., 2024), and reflection (Zhang et al., 2025b, Wang and Liu, 2024); and (2) embedding agentic capabilities directly into LLMs/VLMs to create native agent models (Qin et al., 2025, Guo et al., 2025). Despite these advances, CUAs continue to exhibit significant limitations in their ability to navigate complex computing environments effectively.
These limitations stem from a fundamental challenge illustrated in Figure 1: a profound disconnect between how humans and AI agents perceive and interact with computing environments. In the current paradigm, researches on CUA predominantly focus on enhancing agent capabilities while largely overlooking the environments in which these agents operate. As the right part of Figure 1 illustrates, there exists a significant opportunity to bridge this gap by creating agent-specific environments that mediate between the raw computer interface and the agent’s cognition. Rather than forcing agents to adapt to interfaces designed for human perception and action, we can transform these environments into contextualized representations that align with the semantic understanding of language models.
Inspired by this and based on the foundation laid out by the AIOS architecture (Mei et al., 2024, Rama et al., 2025), we introduce AIOS 1.0, an augmented platform developed to move towards contextualization of computer environments for LLM-based agents. Specifically, AIOS 1.0 aims to model the full computer as MCP [^1] https://modelcontextprotocol.io/introduction server to provide LLM-friendly and well-structured context where interactive elements permissible actions are explicitly encoded in JSON schemas. This contextualization process transforms the computer from an opaque operational environment into a semantic landscape that aligns with the LLM’s understanding of the environment. This design equips agents with an informative yet compact view of the state space, enabling finer-grained reasoning over available operations and facilitating longer-horizon planning. By decoupling interface complexity from decision complexity, AIOS 1.0 addresses a fundamental bottleneck in current CUA capabilities. The contextualized environment acts as a semantic bridge between the agent’s language-based reasoning and the computer’s operational requirements.
Upon AIOS 1.0, we introduce LiteCUA, a computer-use agent that is developed by us with simple architectural design. LiteCUA achieves competitive results with several agent frameworks, which scores 14.66% in the OSWorld benchmark (Xie et al., 2024). The success of LiteCUA emphasizes that AIOS 1.0’s approach of creating agent-specific environments can effectively reduce the implementation complexity required to build computer-use agents, allowing researchers to focus on advancing agent capabilities rather than wrestling with the intricacies of computer interface interpretation. By reframing the computer-use agent challenge as an environmental contextualization problem rather than solely a model capability issue, AIOS 1.0 aims to catalyze the next leap in computer-use agent capabilities. This approach represents a crucial step toward developing truly general computer-use agents, systems that can navigate the digital world with the flexibility and comprehension previously exclusive to human users.
Figure 1: Illustration of the differences between how human operates computer and how computer-use agent (CUA) operates computer, where CUA requires a specific contextualized environment to understand the original computer environment and take actions.
2 Architecture and System Design
Following the architecture of AIOS (Mei et al., 2024, Rama et al., 2025), which abstracts a specific AIOS kernel for isolating agent logic (Rama et al., 2025) and managing agent-related resources and services, we extend this architecture to create AIOS 1.0, a platform optimized for developing and running computer-use agents.
Figure 2: The serving architecture of AIOS 1.0 for computer-use agent, which extends on the basis of AIOS 0.x versions (Mei et al., 2024).
2.1 Architecture
Application Layer: In the Application Layer, agents are provided with SDK APIs to interact with some crucial components like Terminal, Code editor, Browser, and Document. By abstracting the essential interaction patterns common across all interfaces, AIOS 1.0 creates a consistent interaction surface that allows agents to navigate diverse computing environments. This approach significantly reduces the complexity agents face when operating in heterogeneous computer components. The agents can leverage the system-level resources, such as LLM and vector database by revoking the APIs within AIOS SDK (Rama et al., 2025). While maintaining the original design from AIOS SDK regarding LLM, Memory and Storage Query APIs, AIOS 1.0 designs the integration of MCP for tool communications. MCP Client and Http Client are established as communication channels to map the agent’s semantic understanding into computer manipulation commands.
Kernel Layer: AIOS 1.0 extends the AIOS Kernel with significant enhancements focused on computer contextualization. While preserving essential components like LLM Core(s), Context Manager, and Memory Manager, we have fundamentally redesigned the Tool Manager module to incorporate a VM (Virtual Machine) Controller and MCP Server. This redesign creates a sandboxed environment that allows agents to safely interact with computer systems while maintaining a consistent semantic mapping between agent intentions and computer operations. The AIOS system calls are also complimented with Tool related system calls.
2.2 Contextualizing Computers as MCP Servers
The core innovation of AIOS 1.0 lies in its approach to transforming computers from operational tools into contextual environments that language models can naturally comprehend. As shown in Figure 3, our MCP server serves a semantic bridge between agent cognition and computer operation through a structured contextualization process:
Figure 3: Pipeline for agents to take actions to interact with VM environments.
Environment Perception Framework: Our environment perception system implements a multi-modal sensing approach that captures the computer state through complementary channels: Screenshot to provide visual signals and Accessibility Tree to provide structural information within the the current screen. Aside from the information from the visible elements in the screenshot, we also provide the mechanisms to inspect invisible information like the version information of softwares. Through these components, we provide a comprehensive semantic representation of the computing environment. The perception components allows agents to understand the computer environment as a meaningful semantic landscape rather than a collection of pixels or interface elements, benefiting their ability to reason about computer states and potential actions.
Action Space Semantics: Within AIOS 1.0, the action space is constrained into atomic but essential computer-related operations including CLICK, SCROLL, TYPE, DRAG, etc. Each action type is further encoded into the combination of several GUI control signals by leveraging the pyautogui [^2] https://pyautogui.readthedocs.io/en/latest/ library. This action space enables agents to plan and execute complex interaction sequences while reasoning about operations at a conceptual level similar to human thought processes.
VM Controller: Considering that some of the computer-use operations can lead to irreversible outcomes, we design the VM controller to allow agents to reason and act in a safe sandbox. The HTTP interface to the Virtual Machine establishes a standardized communication protocol and serves as the abstraction of low-level operations to allow agents to interact with computer environments regardless of the underlying implementation details.
Through above contextualization, AIOS 1.0 transforms computers from operational challenges into semantic environments that language models can naturally navigate, substantially reducing the cognitive gap between language understanding and computer interaction.
3 LiteCUA: a Computer-Use Agent built on AIOS 1.0
To demonstrate the effectiveness of the AIOS 1.0 platform, we developed LiteCUA, a lightweight computer-use agent that leverages our contextualized environment architecture. We will introduce its design in the following sections.
3.1 Orchestrator-Worker Architecture
As illustrated in Figure 4, LiteCUA adopts an orchestrator-worker architecture that distributes cognitive responsibilities across specialized modules while maintaining cohesive task execution, which is inspired by some popular agent frameworks (Wu et al., 2023, Agashe et al., 2025, Li et al., 2023).
The Orchestrator serves as the central planning and coordination module, responsible for decomposing complex tasks into manageable subtasks, tracking progress toward goal completion. Worker modules perform specialized functions within the agent’s cognitive pipeline. This specialization allows each module to optimize for its specific function without conflating distinct cognitive processes. The modular architecture also facilitates incremental improvements to specific capabilities without requiring comprehensive system redesign.
Figure 4: Design of LiteCUA. We adopt a simple orchestrator-worker architecture, where orchestrator is responsible for planning and assigning tasks and other workers deal with specialized duties like perceiving, reasoning and acting.
3.2 Perceive-Reason-then-Act
In LiteCUA, we deliberately design a perceive-reason-act cognitive cycle that systematically processes computer environments through three specialized stages:
The Perceptor module integrates multi-modal inputs, i.e., screenshots and A11y tree to construct a comprehensive representation of the computing environment. Rather than processing raw inputs directly, the Perceptor organizes perceived elements into semantically meaningful structures that capture both the functional and relational aspects of interface elements. For example, a button is represented not merely as a visual element but as an interactive component with specific usage semantics, e.g., "Type: Button, Usage: Open the chrome".
The Reasoner module serves as the cognitive core of LiteCUA, interpreting the structured environment information in relation to the current task objectives. It processes textual environment description alongside visual screenshot data, applying operational rules that encode interaction patterns and constraints within computing environments. These information will be fed into the Reasoner to generate both thought (interpretative understanding of the current state) and action (strategic planning for next steps) outputs.
The Actor module translates high-level action intentions into precise computer operations through a structured action space. It transforms semantic actions (Type, Scroll, Click, Drag, Wait) to specific GUI control commands leveraging pyautogui, e.g., pyautogui.click(100,100).
This perceive-reason-act cycle enables LiteCUA to maintain goal-directed behavior across extended interaction sequences while adapting to dynamic computing environments. By leveraging AIOS 1.0’s contextualized environment architecture, we demonstrate that effective computer-use agents can be constructed with relatively simple and straightforward architectures and the performance of LiteCUA will be shown in Section 4.
4 Evaluation
4.1 Setup
Deployment Environment.
The experiments are conducted on an isolated VM with 2 AMD Ryzen CPU cores and 4 GB of memory with ubuntu 22.04 desktop.
Benchmark.
We evaluate the LiteCUA computer-use agent built atop of AIOS 1.0 on OSWorld (Xie et al., 2024), a benchmark which provides a diverse online environment for evaluating multi-modal agents on computer-using tasks. It consists of 369 tasks involving real-world web and desktop applications.
Baselines.
For baselines, we choose three LLM baselines, i.e., GPT (4o-mini, GPT-4o) (Achiam et al., 2023), and Gemini-1.5-flash (Team et al., 2023), and three agent baselines which adopt GPT-4o as their backbones, i.e., Friday (Wu et al., 2024), Open-Interpreter (Lucas, 2024), and AgentStore (Jia et al., 2024). For a fair comparison, we evaluate all the methods under the "Screenshot + A11y tree" setting and only compare agents which do not introduce extra models for grounding. The metric we use is success rate provided within OSWorld.
4.2 Comparative Performance
Table 1: Success Rate Comparison
Method (Screenshot + A11y tree) OSWorld
LLMs
GPT-4o-mini (Achiam et al., 2023) 6.21
GPT-4o (Achiam et al., 2023) 11.21
Gemini-1.5-pro (Team et al., 2023) 5.1
Agents (based on GPT-4o)
Friday (Wu et al., 2024) 11.11
Open-Interpreter (Lucas, 2024) 8.94
AgentStore (Jia et al., 2024) 13.55
LiteCUA 14.66
Table 2: Detailed Performance of LiteCUA
Domain Score Average Steps
Chrome 7/46 28.4
GIMP 4/26 32.7
Libreoffice Calc 0/47 50.0
Libreoffice Writer 2/23 42.1
Libreoffice Impress 7/47 31.5
OS 13/24 18.2
VLC 3.11/17 35.8
Thunderbird 0/15 50.0
VSCode 8/23 25.6
Multi-app 11/101 38.4
Total 54.10/369 35.3
From Table 2, LiteCUA achieves a 14.66% success rate on the OSWorld benchmark, showing modest improvements over both standalone model baselines and other agent frameworks. Compared to GPT-4o (11.21%), our approach demonstrates that properly contextualizing computer environments for LLMs can yield better results than using the model in isolation. Among agent frameworks using GPT-4o as their backbone, LiteCUA outperforms Friday (11.11%) and Open-Interpreter (8.94%), with a slight edge over AgentStore (13.55%). These results, while still showing significant room for improvement, suggest that our approach of treating computers as contextual environments provides advantages for computer-use tasks. The AIOS Agent Operating System architecture allows for more effective translation between LLM understanding and computer operations compared to traditional prompt-based approaches. However, the overall modest success rates across all systems (below 15%) indicate that computer-use tasks remain challenging for current AI systems, suggesting that significant architectural and modeling improvements are still needed to achieve robust performance.
4.3 Performance Breakdown by Domains
The domain-specific results reveal clear patterns in LiteCUA’s capabilities and limitations. The system performs best on OS-related tasks (13/24, 54.2%) with the lowest average step count (18.2), likely because these operations involve clearer, more structured interactions. Similarly, VSCode tasks show reasonable performance (8/23, 34.8%) with relatively efficient execution (25.6 steps on average). In contrast, the system struggles significantly with complex productivity applications, achieving no success on Libreoffice Calc (0/47) or Thunderbird (0/15). These domains typically involve more complex screenshots that need to be perceived and have specialized interaction patterns. The maximum step count (50.0) in these domains indicates that the system exhausts its operation budget without achieving the task goals. Multi-application tasks, which require maintaining context across different environments, show modest performance (11/101, 10.9%) with higher average step counts (38.4), reflecting the additional complexity of context switching. Web and media applications (Chrome and VLC) show intermediate performance.
5 Related Work
5.1 Computer-Use Agent
The automation of computer-based tasks through agents has significantly advanced with (M)LLMs, enabling them to operate across desktop, mobile, and web platforms. A key technical improvement involves moving beyond reliance on potentially noisy textual representations like HTML or accessibility trees towards more robust, vision-centric perception. For instance, systems like UGround (Gou et al., 2025) and Aria-UI (Yang et al., 2024b) champion agents that perceive GUIs visually, akin to humans, and perform pixel-level operations. UGround leverages large-scale synthetic web data and adapts the LLaVA architecture for universal visual grounding , while Aria-UI employs a pure-vision approach with a scalable data pipeline to synthesize diverse instruction samples for grounding and incorporates action histories for context-aware reasoning. On mobile platforms, AutoDroid demonstrates a technical advancement by combining LLM commonsense with app-specific knowledge, acquired through automated dynamic analysis and exploration-based memory injection, to automate tasks without prior manual effort (Wen et al., 2024). The evaluation of such computer-use agents is also evolving, with benchmarks like AgentBench (Liu et al., 2023), TheAgentCompany (Xu et al., 2024), and AgentGym (Xi et al., 2023) providing more complex and realistic scenarios for assessing their operational capabilities. Furthermore, the convergence of API-based and GUI-based interaction, creating hybrid agents, is a promising direction for versatile software automation (Qiao et al., 2025).
5.2 AI Agent Operating System
The increasing complexity of LLM-based agents calls for dedicated infrastructure, leading to the development of AI Agent Operating Systems designed to provide foundational services. AIOS (Mei et al., 2024), as the first LLM agent operating system, introduces an LLM-specific kernel to manage resources and LLM-specific services, offering agent scheduling, context management, memory and storage management, tool management, and access control, along with the agent SDK (Rama et al., 2025) for agent development. Such systems address critical challenges like efficient resource allocation, a problem also explored by research into self-resource allocation in multi-agent LLM systems where LLMs themselves can act as planners for task distribution (Amayuelas et al., 2025). UFO2 (Zhang et al., 2025a) designs and implements the desktop agent OS within the Windows environment. These developments show that agent-specific environments are crucial for creating scalable, reliable, and secure ecosystems for advanced agent applications.
6 Conclusion and Envision
In this paper, we introduced AIOS 1.0, a platform that reconceptualizes the computer-use agent challenge by focusing on environmental contextualization rather than model capabilities alone by contextualizing the computer as the MCP server. We also introduce LiteCUA, our demonstrative agent built on the AIOS 1.0 platform, which achieved competitive performance on the OSWorld benchmark with a 14.66% success rate, outperforming several specialized agent frameworks despite its relatively simple architecture.
While our current implementation shows promising results, several directions for future work remain. First, enhancing the perception framework to better capture temporal and causal relationships within computer environments could improve agent performance on complex multi-step tasks. Second, developing more sophisticated action space semantics that incorporate probabilistic reasoning could enable more robust operation in uncertain environments. Finally, extending our approach to additional computing domains, particularly those with specialized interfaces like professional creative applications, could further validate the generalizability of our contextualization methodology. AIOS 1.0 represents a step toward artificial general intelligence systems that can interact with the digital world, not by painstakingly teaching AI to use interfaces designed for humans, but by transforming those interfaces into environments that AI can naturally comprehend. This paradigm shift, from adapting agents to computers toward adapting computers for agents, offers a promising path forward for developing truly general computer-use capabilities in AI systems.